This is a word problem:
An investment fund has $3000 more invested at 8% than it does at 10%. If the annual return from the 8% investment is the same as the annual return from the 10% investment, how much is invested at each rate?
When given the total return value, I can calculate interest. With out it, I'm stuck.

Found 2 solutions by richwmiller, scott8148:
Answer by richwmiller(17219)

Thanks for telling us it is a word problem! i would never have guessed! lol!
y=x+x+3000
.1*x=.08*(x+3000)
x=12000
x+3000=15000

Answer by scott8148(6628)

let x="amount at 10%", so x+3000 is the amount at 8%

(x + 3000)(8%) = (x)(10%)